Let f(x) = 2x - 1 and g(x) = (x + 5)/2. Find the function (f o g)(x). Simplify your answer.
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

Let f(x) = 2x - 1 and g(x) = (x + 5)/2.
Find the function (f o g)(x).
Simplify your answer.
------
fog(x) = f[g(x)] = f[(x+5)/2] = 2[(x+5)/2]-1
---
= (x+5)-1
---
= x+4
===========
Cheers,
Stan H.
===========

Answer by edjones(8007)

f(x) = 2x - 1
g(x) = (x + 5)/2
fog(x)=2((x + 5)/2)-1
=x+5-1
=x+4
.
Ed